OpenStudy (anonymous):

Factor: x2 + 11x + 10 a: (x + 10)(x +1) b: (x + 11)(x + 10) c: (x + 2)(x + 5) d: prime

OpenStudy (usukidoll):

To factor \[x^2+11x+10 \] we have to focus on the middle term and the last term so our last term is 10, and we have to list all possibilities of multiplication of 10 so our choices would be 1 x 10 = 10 2 x 5 = 10 then based on the middle term we use addition and only one pair of numbers work

OpenStudy (anonymous):

ok

OpenStudy (usukidoll):

so what's 1+10? what's 2+5 ?

OpenStudy (anonymous):

11 and 7

OpenStudy (anonymous):

prime?

OpenStudy (usukidoll):

ah! We found our 11 because 10+1 = 11 10 x 1 = 10 so since we have found or numbers we can factor

OpenStudy (anonymous):

ok

OpenStudy (usukidoll):

since your quadratic equation is all positive signs, we don't have to worry about sign combinations

OpenStudy (usukidoll):

so we have (x+ ?)(x + ? ) fill in those ? using the information I've typed earlier

OpenStudy (anonymous):

x+11

OpenStudy (usukidoll):

close xD "We found our 11 because 10+1 = 11"

OpenStudy (anonymous):

x+10?

OpenStudy (usukidoll):

no.. remember those 2 digital 10 and 1 that produced 11? write those down (x+?)(x+?)

OpenStudy (usukidoll):

wut did I just type .. it's 2 numbers 10 and 1 =_=

OpenStudy (anonymous):

5+2

OpenStudy (usukidoll):

no now we're going backwards 5+2 is 7 and it doesn't match the middle term which is 11 so we throw that combination out

OpenStudy (anonymous):

then i don't know

OpenStudy (usukidoll):

:/ we found our combination 10 x 1 = 10 which matches the last term 10 + 1 = 11 matches the middle term what numbers did I just use to make 10 and 11 ?

OpenStudy (anonymous):

1

OpenStudy (usukidoll):

yeah that was one of the numbers. (x+1)(x+?) what was the other number?

OpenStudy (anonymous):

was it 2?

OpenStudy (usukidoll):

no

OpenStudy (anonymous):

and 10

OpenStudy (usukidoll):

yes! (x+1)(x+10)

OpenStudy (anonymous):

thanks

OpenStudy (usukidoll):

and you can check by using the FOIL method (x+1)(x+10) = x^2+10x+x+10 =x^2+11x+10

OpenStudy (anonymous):

so is that the final answer?

OpenStudy (usukidoll):

yeah